OpenStudy (anonymous):

Choose the correct simplification of 5^4\5^2

OpenStudy (anonymous):

i belive it is 25

OpenStudy (unklerhaukus):

\[\frac{5^4}{5^2}=\frac{5\times 5\times 5\times 5\times 5}{5\times 5}\]

OpenStudy (unklerhaukus):

ops one too many fives on the numerator, but you get the point two fives cancel and your left with two ontop

OpenStudy (anonymous):

\[{{5^4}\over{5^2}}={5^{4-2}} = {5^2}\]